# LLM-based Agents Suffer from Hallucinations: A Survey of Taxonomy, Methods, and Directions

### v1

@@ 4. Systematic Methodology
In this section, we provide a comprehensive review of recent methods for mitigating and detecting hallucinations in agents. For mitigation, we systematically group existing approaches into three main categories in Section 4.1: (1) Knowledge Utilization, which leverages external and internal knowledge to address hallucination issues; (2) Paradigm Improvement, which focuses on developing advanced learning paradigms to prevent agent hallucinations during training and inference; and (3) Post-hoc Verification, which introduces verification techniques to calibrate agent outputs and reduce the risk of hallucinations. In Fig. 4, we provide a simple illustration of each specific mitigation approach. For detection, existing methods are fewer compared to mitigation, so we place detection methods after mitigation methods. In Section 4.2, we review the available detection approaches corresponding to each type of agent hallucinations.
4.1. Agent Hallucination Mitigation
4.1.1. Knowledge Utilization
@@
Knowledge utilization equips LLM-based agents with accurate and reliable information support, enabling multi-dimensional regulation of their reasoning, execution, perception, and memorization behaviors. This, in turn, helps reduce hallucinatory outputs stemming from knowledge gaps and biases. Here we categorize knowledge utilization into two types: (1) External Knowledge Guidance: Agent hallucinations often arise from insufficient knowledge guidance. Therefore, external knowledge serves as a critical resource for mitigating hallucinations and enhancing output reliability. (2) Internal Knowledge Enhancement: This line of research focuses on activating or rectifying the agent’s internal knowledge to better mitigate hallucinations.
External Knowledge Guidance
@@
We categorize the existing approaches of external knowledge guidance into two types: Expert Knowledge and World Models. Expert knowledge refers to the domain-specific information distilled from the behaviors and decision-making processes of human experts or high-performing systems during task execution. Owing to its high reliability and precision, expert knowledge serves as a robust external reference, significantly reducing the likelihood of hallucinations across various agentic operations. In related literature, expert knowledge can be manifested different forms, and we list several representative forms: (1) Knowledge base methods: Expert knowledge can be stored as knowledge bases, such as introspective reasoning knowledge bases (Liang et al., 2024), action knowledge bases (Zhu et al., 2024), experiential knowledge bases (Fu and Bian, 2025), from which agents can retrieve real cases and incorporate them into prompts to support more accurate decision-making (Wang et al., 2025g; Hatalis et al., 2025; Li et al., 2025b). (2) Rule-based methods: expert knowledge can be leveraged to establish rules such as mathematical rules (Tudor et al., 2025; Liu et al., 2025a; Wang and Casella, [n. d.]), linguistic rules (Chu et al., 2025; Lu et al., 2025; Ni et al., 2025), procedural rules (Lumer et al., 2025a; Yuan et al., 2024b; Wu et al., 2024c) or MAS communication protocols (Chen et al., 2024i; Huang et al., 2025a; Kwon and Zhang, 2025). These rules can explicitly constrain the agent’s external behaviors, mitigating agent hallucination risks (Wang et al., 2025h; Zhao et al., 2025a; Hsieh et al., 2023). (3) Data construction methods: By leveraging expert annotations and prompting techniques, synthetic datasets can include diverse and targeted instruction–response pairs (Abdelaziz et al., 2024; He et al., 2025b; Liu et al., 2023b), Q&A pairs (Jain et al., 2024; Wang et al., 2024e; Dong et al., 2025) or preference pairs (Lou et al., 2024; Chen et al., 2024j; Huang et al., 2025b). These curated datasets can enhance the agent’s capability to handle complex scenarios, including fine-grained reasoning, tool usage, and multi-modal perception.
@@
World models refer to systems that construct useful representations of the world, equipping agents with foundational world knowledge (Long et al., 2025; Zhang et al., 2025c; Zhao et al., 2025c). This enables them to handle real-world problems and reject outputs that contradict established facts, thereby effectively mitigating the occurrence of agent hallucinations (Richens et al., 2025; Burchi and Timofte, [n. d.]). In POMDP setting, world models are particularly important for constructing more accurate estimations of partially observable environments (Hu and Shu, 2023; Zhang et al., 2024h). The scope of world knowledge encompasses a wide range of domains, from fundamental Physical Principles (Duan and Wang, 2024; Zhang et al., 2024c; Li et al., 2025e) such as Newton’s laws of motion and the conservation of mass to complex Social Systems and Structures (Zhao et al., 2023a; Chae et al., 2024). For example, NavMorph (Yao et al., 2025) leverages the world model to provide knowledge of the physical environment, avoiding navigation biases caused by environmental unfamiliarity, while WKM (Qiao et al., 2024) supplies common-sense knowledge—such as the need to clean ingredients before cooking—to achieve behaviors aligned with human expectations.
Internal Knowledge Enhancement
@@
Current enhancement strategies can be broadly classified into two types: Internal Knowledge Activation and Internal Knowledge Rectification. The former aims to leverage Prompt Engineering to carefully stimulate and utilize the agent’s internal knowledge, guiding it to fully exploit its capabilities. Here we highlight several representative prompting techniques: (1) Chain-of-Thought (CoT): By incorporating step-by-step reasoning instructions (Wei et al., 2022b; Manas et al., 2024) into the prompt, CoT guides the agent to break down complex problems and output intermediate steps in a logical sequence, thereby enhancing the accuracy and reliability of handling multi-step tasks. (2) Tree-of-Thought (ToT): Building upon CoT, ToT explores and evaluates multiple reasoning paths in parallel (Yao et al., 2023a; Wang et al., 2022b; Chu et al., 2024a), fully utilizing the diversified information within the agent’s internal knowledge to strengthen the completeness and robustness of its reasoning process. (3) Constrained Prompting: By incorporating various forms of constraints (Singh et al., 2023; Goncharov and Dansereau, 2025) such as semantic (Chen et al., 2024e) and spatial constraints (You et al., 2023; Wu et al., 2025c) into prompts, this approach guides the agent to focus on task-relevant information, effectively reducing the generation of irrelevant content.
@@
Internal knowledge rectification generally involves two representative paradigms: (1) Knowledge Editing replaces inaccurate or outdated knowledge with correct knowledge while minimizing disruptions to other internal knowledge (Wang et al., 2024d; Li et al., 2024b; Fang et al., 2025; Gu et al., 2024). It is typically divided into two main approaches (Yao et al., [n. d.]): the locate-and-edit approach (Li et al., 2023c; Dai et al., 2022), which identifies and modifies specific knowledge components; the meta-learning approach (De Cao et al., 2021; Mitchell et al., 2022), which aims to to learn how to update the model’s parameters for adjusting knowledge for new information. (2) Knowledge Unlearning, in contrast to knowledge editing, focuses on removing erroneous knowledge to improve the agent’s reliability. It includes two main approaches (Si et al., 2023): parameter optimization (Sun et al., 2025b; Yao et al., 2024; Eldan and Russinovich, 2023; Yoon et al., 2025a), which fine-tunes model parameters under certain conditions (e.g., parameter update scopes(Chen and Yang, [n. d.]; Wu et al., 2023)) to forget some targeted knowledge; parameter merging (Ilharco et al., [n. d.]; Zhang et al., 2023a), which performs offline operations on model parameters without further training.
@@
Figure 4. A simple illustration of approaches to agent hallucination mitigation. It encompasses three branches, knowledge utilization, paradigm improvement, and post-hoc certification, comprising a total of ten representative methods.
4.1.2. Paradigm Improvement
@@
This section focuses on enhancing the agent’s capabilities through general learning paradigms to mitigate hallucinations. We summarize six representative paradigms as follows. Contrastive Learning is a well-established self-supervised learning paradigm whose core idea is to learn more discriminative representations by comparing similarities and differences between samples (Le-Khac et al., 2020; Cao et al., 2021; Lin et al., 2025a). By incorporating contrastive learning, the agent can better identify relevant and accurate patterns, reducing errors when handling unfamiliar or incomplete task inputs and effectively mitigating hallucination issues (Chia et al., 2023; Li et al., 2025a; Zhao et al., 2023b; Jiang et al., 2024).
@@
Curriculum Learning is a learning paradigm inspired by the way humans learn. Its core idea is to start training the model with ”easy” examples and gradually move to ”harder” ones, improving learning efficiency and generalization ability (Le-Khac et al., 2020). By leveraging curriculum learning, agents rapidly accumulate successful experiences by starting with simpler tasks and gradually transferring this knowledge to more complex scenarios. This progressive learning paradigm enhances their fundamental capabilities, effectively mitigating agent hallucinations (Deng et al., 2024b; Xi et al., 2024; He et al., 2025b; Gao et al., 2024).
@@
Reinforcement Learning serves as a fundamental learning paradigm, emulating the biological learning process of ”trial, feedback, and adjustment”. This approach enables agents to interact with their environment and learn to maximize cumulative rewards through iterative trial and error (Kaelbling et al., 1996). Through the reinforcement learning paradigm, the agent can adjust its policy based on rewards, thereby optimizing future decision-making, learning to achieve goals in complex environments, and significantly mitigating potential hallucination issues in multi-turn interactions (Xiong et al., 2025b; Shen et al., 2025; Xu et al., 2024a; Ben-Kish et al., 2023; Yu et al., 2024b; Xu et al., 2025a). In addition, there are also studies that aim to improve the reliability of reward signals in order to enhance the overall model effectiveness (Zhang et al., 2024f). Causal Learning refers to the process in statistical machine learning that goes beyond capturing mere correlations and instead aims to model, discover, and leverage causal relationships between variables, thereby enhancing the model’s interpretability and generalization ability (Pearl, 2009; Lin et al., 2025b). In the context of LLM-based agents, identifying and leveraging potential causal relationships within task inputs and decision-making processes enables more accurate understanding of how different factors influence task outcomes. This helps reduce representational errors and logical inconsistencies for mitigating perception (Li et al., 2025c; Hu et al., 2025a) and reasoning hallucinations (Wang et al., 2024i; Tang et al., 2023b).
@@
Table 1. An overview of methods for mitigating various types of agent hallucinations. To save space, here we abbreviate hallucination types and mitigation methods mentioned earlier. ”✔” indicates that related works already exist; ”–” indicates that it has not yet been explored.
@@
Graph Learning is a class of machine learning paradigms specifically designed to handle graph-structured data, which has been widely applied in various domains such as social networks, bioinformatics, and recommendation systems (Xia et al., 2021; Lin et al., 2024; Zhu et al., 2025b). It enhances the agent’s capability to organize and manage tasks in a structured and systematic manner, particularly for tool usage and memory updates, thereby reducing the risk of agent hallucinations (Bei et al., 2025; Wu et al., 2024a; Anokhin et al., 2024; Liu et al., 2025h). For example, ControlLLM (Liu et al., 2023a) builds a tool graph comprising resource and tool nodes, and employs a parallel graph search algorithm to efficiently find appropriate tool paths for mitigating execution hallucinations. More importantly, recent studies leverage graph learning to optimize communication topology in LLM-based MAS (Zhuge et al., 2024; Feng et al., 2025b; Yang et al., 2025b; Chen et al., 2024i). By dynamically learning which agents should communicate, when, and through what channels, the system can suppress noisy message passing while promoting informative interactions.
@@
Unlike the aforementioned training-phase learning paradigms, Decoding Optimization is a test-time learning paradigm. This paradigm can improve output quality by adjusting probability distributions or attention patterns in decoding processes, ensuring better alignment with inputs and factual knowledge,thereby reducing hallucinations stemming from reasoning errors (Chuang et al., 2023; Shi et al., 2024a; Huang et al., 2024a; Wang et al., 2024b; Tang et al., 2025b). Common decoding strategies include contrastive decoding, self-calibrated attention decoding (Tang et al., 2025a) and fusion decoding (Wang et al., 2024b). Contrastive decoding (Leng et al., 2024; Park et al., 2025) selects more reliable and context-consistent outputs by comparing multiple decoding candidates. Self-calibrated attention decoding (Tang et al., 2025a) aims to dynamically adjust the attention weights to capture more informative cues. Fusion decoding incorporates the multi-source information to reduce dependence on single-modal priors for mitigating hallucinations. CGD (Deng et al., 2024a) is one of representative works, which integrates semantic similarity scores between textual and visual content with final-layer token likelihoods to reduce reliance on language priors.
4.1.3. Post-hoc Verification
@@
This paradigm focuses on monitoring and evaluating outputs after task execution. It generally follows a multi-step process in which, immediately after each step, the agent assesses the validity, consistency, and factuality of the intermediate decisions and actions generated in that specific step. This assessment helps prevent the accumulation and propagation of hallucinations in long-horizon tasks. Existing approaches can be broadly categorized into two types: (1) Self-verification Mechanism, in which the agent introspectively reviews its own behaviors using an internal reasoning strategy or prompt-based self-assessment. (2) Validator Assistance, which leverages another independent system to detect potential flaws in the agent’s outputs.
Self-verification Mechanism
@@
Self-verification is a lightweight and model-internal approach wherein agents assess the validity and reliability of their own outputs without relying on external validators. It plays a crucial role in hallucination mitigation (Moncada-Ramirez et al., 2025), and several technical methods have been explored to implement self-verification. Specifically, Self-reflection enables agents to revisit and critique their own outputs, often through prompting techniques that encourage introspection and identification of reasoning flaws (Qu et al., 2025a; Lee et al., 2024; Shinn et al., 2023; Rasmussen et al., 2025), which can be further facilitated by estimating the agent’s own confidence or uncertainty (Bhatt et al., 2025; Zhang et al., 2024e; Hu et al., 2024b; Chen et al., 2025c; Zhang et al., 2025b). Self-consistency leverages the generation of multiple candidate outputs, such as diverse reasoning paths or answers, and aggregates them using majority voting or confidence-weighted schemes to select the most reliable results (Wang et al., 2022b; Liang et al., 2023; Wu and Cardie, 2025). Self-questioning guides the agent to pose and answer critical verification questions grounded in its own reasoning process, enabling the detection of unsupported assertions (Karov et al., 2025; Ji et al., 2023b; Manakul et al., 2023). Collectively, these strategies establish a foundation for autonomous error detection and correction, empowering agents to self-regulate hallucination risks.
@@
Figure 5. A typology of detection methods, highlighting representative approaches for each type of agent hallucination.
Validator Assistance
This approach leverages external validators to verify the correctness of an agent’s outputs, aiming to mitigate hallucinations (Cemri et al., 2025; Zhang et al., 2025g). According to the types of external validators, we divide existing methods into the following five categories: Language-based Validators independently assess the truthfulness or coherence of an agent’s outputs using techniques such as atomic fact decomposition and entailment checking (Jalaian et al., 2025; Wan et al., 2024a; Xue et al., 2025). Retrieval-based Validators rely on some external sources such as search engines to verify whether outputs aligns with established facts (Belyi et al., 2024; Gou et al., 2024). Execution-based Validators evaluate outputs by running generated codes or plans in external execution environments, enabling direct assessment of correctness through functional outcomes (Ni et al., 2023; Zhang et al., 2023c). Simulation-based Validators validate agent behaviors through interaction with sandboxed environments, allowing for realistic testing in tasks involving embodiment, planning, or sequential control (Kleiman et al., 2025). Building on these categories, an increasing number of studies have explored Ensemble-based Validators that integrate multiple types of validators to improve robustness. By enabling cross-verification among different approaches, these methods help mitigate the limitations inherent in individual validation strategies (Wu and Cardie, 2025). In the end, Table 1 summarizes which types of agent hallucinations can be addressed by the ten listed mitigation methods.
4.2. Agent Hallucination Detection

### v2

@@ IV Systematic Methodology
In this section, we provide a comprehensive review of recent methods for mitigating and detecting hallucinations in agents. For mitigation, we systematically group existing approaches into three main categories in Section IV-A: 1) Knowledge Utilization, which leverages external and internal knowledge to address hallucination issues; 2) Paradigm Improvement, which focuses on developing advanced learning paradigms to prevent agent hallucinations during training and inference; and 3) Post-hoc Verification, which introduces verification techniques to calibrate agent outputs and reduce the risk of hallucinations. In Fig. 3, we provide a simple illustration of each specific mitigation approach. For detection, existing methods are fewer compared to mitigation, so we place detection methods after mitigation methods. In Section IV-B, we review the available detection approaches corresponding to each type of agent hallucinations.
IV-A Agent Hallucination Mitigation
@@
1) Knowledge Utilization. This category equips LLM-based agents with accurate and reliable information support, enabling multi-dimensional regulation of their reasoning, execution, perception, and memorization behaviors. This, in turn, helps reduce hallucinatory outputs stemming from knowledge gaps and biases. Here we categorize knowledge utilization into two types: a) External Knowledge Guidance: Agent hallucinations often arise from insufficient knowledge guidance. Therefore, external knowledge serves as a critical resource for mitigating hallucinations and enhancing output reliability. b) Internal Knowledge Enhancement: This line of research focuses on activating or rectifying the agent’s internal knowledge to better mitigate hallucinations.
a) External Knowledge Guidance (EKG) [^4] The abbreviation is added here mainly to save space in Table I.. We categorize the existing approaches of external knowledge guidance into two types: Expert Knowledge and World Models.
@@
Expert knowledge refers to the domain-specific information distilled from the behaviors and decision-making processes of human experts or high-performing systems during task execution. Owing to its high reliability and precision, expert knowledge serves as a robust external reference, significantly reducing the likelihood of hallucinations across various agentic operations. In related literature, expert knowledge can be manifested different forms, and we list several representative forms: i) Knowledge base methods: Expert knowledge can be stored as knowledge bases, such as introspective reasoning knowledge bases [liang2024introspective], action knowledge bases [zhu2024knowagent], experiential knowledge bases [fu2025knowmap], from which agents can retrieve real cases and incorporate them into prompts to support more accurate decision-making [wang2025medagent]. ii) Rule-based methods: expert knowledge can be leveraged to establish rules such as mathematical rules [liu2025enhancing], linguistic rules [chu2025llm+], procedural rules [yuan2024easytool] or MAS communication protocols [kwon2025cp]. These rules can explicitly constrain the agent’s external behaviors, mitigating agent hallucination risks [hsieh2023tool]. iii) Data construction methods: By leveraging expert annotations and prompting techniques, synthetic datasets can include diverse and targeted instruction–response pairs [abdelaziz2024granite], Q&A pairs [dong2025insight] or preference pairs [huang2025enhance]. These curated datasets can enhance the agent’s capability to handle complex scenarios, including fine-grained reasoning, tool usage, and multi-modal perception.
@@
World models refer to systems that construct useful representations of the world, equipping agents with foundational world knowledge [long2025survey]. This enables them to handle real-world problems and reject outputs that contradict established facts, thereby effectively mitigating the occurrence of agent hallucinations [richens2025general]. In POMDP setting, world models are particularly important for constructing more accurate estimations of partially observable environments [hu2023language, zhang2024combo]. The scope of world knowledge encompasses a wide range of domains, from fundamental Physical Principles [duan2024enhancing] such as Newton’s laws of motion and the conservation of mass to complex Social Systems and Structures [zhao2023large, chae2024web]. For example, NavMorph [yao2025navmorph] leverages the world model to provide knowledge of the physical environment, avoiding navigation biases caused by environmental unfamiliarity, while WKM [qiao2024agent] supplies common-sense knowledge—such as the need to clean ingredients before cooking—to achieve behaviors aligned with human expectations.
@@
b) Internal Knowledge Enhancement (IKE). Current enhancement strategies can be broadly classified into two types: Internal Knowledge Activation and Internal Knowledge Rectification. The former aims to leverage Prompt Engineering to carefully stimulate and utilize the agent’s internal knowledge, guiding it to fully exploit its capabilities. Here we highlight several representative prompting techniques: i) Chain-of-Thought (CoT): By incorporating step-by-step reasoning instructions [wei2022chain, manas2024cot] into the prompt, CoT guides the agent to break down complex problems and output intermediate steps in a logical sequence, thereby enhancing the accuracy and reliability of handling multi-step tasks. ii) Tree-of-Thought (ToT): Building upon CoT, ToT explores and evaluates multiple reasoning paths in parallel [wang2022self, chu2024beamaggr], fully utilizing the diversified information within the agent’s internal knowledge to strengthen the completeness and robustness of its reasoning process. iii) Constrained Prompting: By incorporating various forms of constraints [singh2023progprompt, goncharov2025segment] such as semantic [chen2024refining] and spatial constraints [wu2025mitigating] into prompts, this approach guides the agent to focus on task-relevant information, effectively reducing the generation of irrelevant content.
@@
Internal knowledge rectification generally involves two representative paradigms. The first is Knowledge Editing which replaces inaccurate or outdated knowledge with correct knowledge while minimizing disruptions to other internal knowledge [li2024banishing]. It is typically divided into two main approaches [yao2023editing]: the locate-and-edit approach [Li2023PMETPM], which identifies and modifies specific knowledge components; the meta-learning approach [de2021editing], which aims to to learn how to update the model’s parameters for adjusting knowledge for new information. The second is Knowledge Unlearning. In contrast to knowledge editing, it focuses on removing erroneous knowledge to improve the agent’s reliability. It includes two main approaches [si2023knowledge]: parameter optimization [sun2025unlearning], which fine-tunes model parameters under certain conditions (e.g., parameter update scopes[chen2023unlearn]) to forget some targeted knowledge; parameter merging [ilharcoediting], which performs offline operations on model parameters without further training.
@@
Figure 3: A simple illustration of approaches to agent hallucination mitigation. It encompasses three branches, knowledge utilization, paradigm improvement, and post-hoc certification, comprising a total of ten representative methods.
2) Paradigm Improvement. This section focuses on enhancing the agent’s capabilities through general learning paradigms to mitigate hallucinations. We summarize six representative paradigms as follows.
@@
a) Contrastive Learning (CoL) is a well-established self-supervised learning paradigm whose core idea is to learn more discriminative representations by comparing similarities and differences between samples [le2020contrastive, cao2021bipartite, lin2025contrastive]. By incorporating contrastive learning, the agent can better identify relevant and accurate patterns, reducing errors when handling unfamiliar or incomplete task inputs and effectively mitigating hallucination issues [chia2023contrastive, li2025refine, jiang2024hallucination].
@@
b) Curriculum Learning (CuL) is a learning paradigm inspired by the way humans learn. Its core idea is to start training the model with ”easy” examples and gradually move to ”harder” ones, improving learning efficiency and generalization ability [le2020contrastive]. By leveraging curriculum learning, agents rapidly accumulate successful experiences by starting with simpler tasks and gradually transferring this knowledge to more complex scenarios. This progressive learning paradigm enhances their fundamental capabilities, effectively mitigating agent hallucinations [deng2024can, xi2024training, r33].
@@
c) Reinforcement Learning (RL) serves as a fundamental learning paradigm, emulating the biological learning process of ”trial, feedback, and adjustment”. This approach enables agents to interact with their environment and learn to maximize cumulative rewards through iterative trial and error [kaelbling1996reinforcement]. Through the reinforcement learning paradigm, the agent can adjust its policy based on rewards, thereby optimizing future decision-making, learning to achieve goals in complex environments, and significantly mitigating potential hallucination issues in multi-turn interactions [xiong2025mpo, shen2025satori, r43, ben2023mocha, yu2024rlhf]. In addition, there are also studies that aim to improve the reliability of reward signals in order to enhance the overall model effectiveness [zhang2024longreward].
d) Causal Learning (CaL) refers to the process in statistical machine learning that goes beyond capturing mere correlations and instead aims to model, discover, and leverage causal relationships between variables, thereby enhancing the model’s interpretability and generalization ability [pearl2009causal, lin2025generative]. In the context of LLM-based agents, identifying and leveraging potential causal relationships within task inputs and decision-making processes enables more accurate understanding of how different factors influence task outcomes. This helps reduce representational errors and logical inconsistencies for mitigating perception [li2025treble, hu2025causal] and reasoning hallucinations [wang2024csce, tang2023towards].
@@
TABLE I: An overview of mitigation methods. To save space, we use the abbreviations of hallucination types and mitigation methods mentioned earlier. “✔” indicates that related works already exist; “–” indicates that it has not yet been explored.
@@
e) Graph Learning (GL) is a class of machine learning paradigms specifically designed to handle graph-structured data, which has been widely applied in various domains such as social networks, bioinformatics, and recommendation systems [xia2021graph, lin2024graph, zhu2025ttgl]. It enhances the agent’s capability to organize and manage tasks in a structured and systematic manner, particularly for tool usage and memory updates, thereby reducing the risk of agent hallucinations [bei2025graphs, wu2024graph, liu2025graph]. For example, ControlLLM [2023controlllm] builds a tool graph comprising resource and tool nodes, and employs a parallel graph search algorithm to efficiently find appropriate tool paths for mitigating execution hallucinations. More importantly, recent studies leverage graph learning to optimize communication topology in LLM-based MAS [zhuge2024gptswarm, yang2025agentnet, chen2024internet]. By dynamically learning which agents should communicate, when, and through what channels, the system can suppress noisy message passing while promoting informative interactions.
@@
Unlike the aforementioned training-phase learning paradigms, f) Decoding Optimization (DO) is a test-time learning paradigm. This paradigm can improve output quality by adjusting probability distributions or attention patterns in decoding processes, ensuring better alignment with inputs and factual knowledge,thereby reducing hallucinations stemming from reasoning errors [huang2024opera, wang2024mllm, tang2025seeing]. Common decoding strategies include contrastive decoding, self-calibrated attention decoding [tang2025intervening] and fusion decoding [wang2024mllm]. Contrastive decoding [leng2024mitigating] selects more reliable and context-consistent outputs by comparing multiple decoding candidates. Self-calibrated attention decoding [tang2025intervening] aims to dynamically adjust the attention weights to capture more informative cues. Fusion decoding incorporates the multi-source information to reduce dependence on single-modal priors for mitigating hallucinations. CGD [deng2024seeing] is one of representative works, which integrates semantic similarity scores between textual and visual content with final-layer token likelihoods to reduce reliance on language priors.
@@
3) Post-hoc Verification. This paradigm focuses on monitoring and evaluating outputs after task execution. It generally follows a multi-step process in which, immediately after each step, the agent assesses the validity, consistency, and factuality of the intermediate decisions and actions generated in that specific step. This assessment helps prevent the accumulation and propagation of hallucinations in long-horizon tasks. Existing approaches can be broadly categorized into two types: a) Self-verification Mechanism, in which the agent introspectively reviews its own behaviors using an internal reasoning strategy or prompt-based self-assessment. b) Validator Assistance, which leverages another independent system to detect potential flaws in the agent’s outputs.
@@
a) Self-verification Mechanism (SM). Self-verification is a lightweight and model-internal approach wherein agents assess the validity and reliability of their own outputs without relying on external validators. It plays a crucial role in hallucination mitigation [moncada2025agentic], and several technical methods have been explored to implement self-verification. Specifically, Self-reflection enables agents to revisit and critique their own outputs, often through prompting techniques that encourage introspection and identification of reasoning flaws [qu2025from, lee-etal-2024-volcano, shinn2023reflexion], which can be further facilitated by estimating the agent’s own confidence or uncertainty [bhatt2025know, hu2024uncertainty, chen2025learning]. Self-consistency leverages the generation of multiple candidate outputs, such as diverse reasoning paths or answers, and aggregates them using majority voting or confidence-weighted schemes to select the most reliable results [wang2022self, liang2023encouraging, wu2025reasoning]. Self-questioning guides the agent to pose and answer critical verification questions grounded in its own reasoning process, enabling the detection of unsupported assertions [karov2025attentive, ji-etal-2023-towards, manakul2023selfcheckgpt]. Collectively, these strategies establish a foundation for autonomous error detection and correction, empowering agents to self-regulate hallucination risks.
b) Validator Assistance (VA). This approach leverages external validators to verify the correctness of an agent’s outputs, aiming to mitigate hallucinations [cemri2025multi, zhang2025agent]. According to the types of external validators, we divide existing methods into the following five categories: Language-based Validators independently assess the truthfulness or coherence of an agent’s outputs using techniques such as atomic fact decomposition and entailment checking [jalaian2025hydra, wan-etal-2024-knowledge, xue2025verify]. Retrieval-based Validators rely on some external sources such as search engines to verify whether outputs aligns with established facts [belyi2024luna]. Execution-based Validators evaluate outputs by running generated codes or plans in external execution environments, enabling direct assessment of correctness through functional outcomes [ni2023lever]. Simulation-based Validators validate agent behaviors through interaction with sandboxed environments, allowing for realistic testing in tasks involving embodiment, planning, or sequential control [kleiman2025simulation]. Building on these categories, an increasing number of studies have explored Ensemble-based Validators that integrate multiple types of validators to improve robustness. By enabling cross-verification among different approaches, these methods help mitigate the limitations inherent in individual validation strategies [wu2025reasoning]. In the end, Table I summarizes which types of agent hallucinations can be addressed by the ten listed mitigation methods.
@@
Figure 4: A typology of methods of agent hallucination detection. We highlight the representative approaches for each type of agent hallucinations.
IV-B Agent Hallucination Detection